OpenStudy (anonymous):

Which of the following equations represents a line which is perpendicular to the line that passes through the points below? (-2,-6) (0,-2) y=-2x+2 y=1/2x +2 y=-1/2x +2 y=2x+2

OpenStudy (anonymous):

if the outcome is -1 then it perpendicular

OpenStudy (anonymous):

i do not know what to choose

OpenStudy (anonymous):

you have to solve it

OpenStudy (anonymous):

y=-2x+2 y=1/2x +2 y=-1/2x +2 y=2x+2